OpenStudy (anonymous):

what is the slope between (2,3)(6,4)

OpenStudy (callisto):

\[slope = \frac{y_2-y_1}{x_2-x_1}\]\[(x_1, y_1) = (2,3)\]\[(x_2, y_2) = (6,4)\]

OpenStudy (anonymous):

so whats the asnwer?

OpenStudy (luigi0210):

Plug in the values

OpenStudy (anonymous):

just gimme the answer please?

OpenStudy (luigi0210):

(4-3)/(6-2) =?